OpenStudy (anonymous):

Multiply. Express your answer in lowest terms. 1/3 times 2/9 = ?

OpenStudy (anonymous):

metal if helped !!!!!!!!!!!!!

OpenStudy (anonymous):

is any body there

OpenStudy (anonymous):

mathmate

OpenStudy (anonymous):

mathmate???

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

hay

OpenStudy (anonymous):

can u answer my question

OpenStudy (mathmate):

\(\frac{1}{3}\times\frac{2}{9}=\frac{1\times3}{2\times9}=\frac{3}{18}=\frac{1}{6}\)

OpenStudy (anonymous):

thnks sooooooooooo much

OpenStudy (anonymous):

so the answer is 1/6

OpenStudy (mathmate):

Wait, there's a typo. \frac{1}{3}\times\frac{3}{9}=\frac{1\times3}{3\times9}=\frac{3}{27}=?

OpenStudy (mathmate):

\(\frac{1}{3}\times\frac{3}{9}=\frac{1\times3}{3\times9}=\frac{3}{27}=\) ?

OpenStudy (mathmate):

Still not right: \(\frac{1}{3}\times\frac{2}{9}=\frac{1\times2}{3\times9}=\frac{2}{27}\)